OpenStudy (anonymous):

Express the following function, F(x) as a composition of two functions f and g. f(x)= x^2/(x^2+4) @misty1212

OpenStudy (misty1212):

you have many choices here, but the easiest one is probably \(g(x)=x^2\) and \(f(x)=\frac{x}{x+4}\)

OpenStudy (misty1212):

then if you compose them you get \[f(g(x))=f(x^2)=\frac{x^2}{x^2+4}\]

OpenStudy (anonymous):

wait can you show me how you got that?

OpenStudy (misty1212):

how i got which part?

OpenStudy (anonymous):

the answer

OpenStudy (misty1212):

\[f(g(x))=f(x^2)=\frac{x^2}{x^2+4}\] this?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

that is how you compose functions if \(g(x)=x^2\) then \[f(g(x))=f(x^2)\]

OpenStudy (misty1212):

this is kind of a crappy explanation, lets see if i can do better

OpenStudy (anonymous):

I'm a bit confused....which is the answer?

OpenStudy (misty1212):

\[f(g(x))=\frac{x^2}{x^2+4}\] is the question your job is to come up with an \(f\) and a \(g\) that work

OpenStudy (misty1212):

when you see \(\frac{x^2}{x^2+4}\) the first thing you notice is that the variable is squared top and bottom right?

OpenStudy (anonymous):

right

OpenStudy (misty1212):

that is why i picked the "inside function " \(g(x)\) as \(g(x)=x^2\)

OpenStudy (misty1212):

then the outside part, since we already know the variable is going to be square, looks something like \[f(\spadesuit)=\frac{\spadesuit}{\spadesuit+4}\]

OpenStudy (misty1212):

if you replace \(\spadesuit\) by \(x^2\) you get what you want\[\frac{x^2}{x^2+4}\]

OpenStudy (misty1212):

so make the inside function \(g(x)=x^2\) and the outside function \(f(x)=\frac{x}{x+4}\) that way you get what you want when you compose them